OpenStudy (anonymous):

the mass of pluto is about 1.31x10^19 metric tons. The mass of the sun is about 1.998x10^27 metric tons. About how many times the mass of pluto is the mass of the sun? give answer in scientific notation

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

divide the mass of the sun by the mass of pluto. that will tell you how much more massive the sun is in relation to pluto

OpenStudy (anonymous):

yes divide

OpenStudy (anonymous):

\[\frac{1.998\times 10^{27}}{1.31\times 10^{19}}\] \[\frac{1.998}{1.31}\times 10^{27-19}\] \[\frac{1.998}{1.31}\times 10^8\]

OpenStudy (anonymous):

lol if i knew how to do it I would

OpenStudy (anonymous):

the first number you get with a calculator.

OpenStudy (anonymous):

i get 1.525 rounded so answer is \[1.525\times 10^8\] steps clear?